OpenStudy (anonymous):

Calculate the double integral ∫∫xcos(2x+y)dA where R is the region 0<=x<=2pi/6, 0<=y<=2pi/4.

OpenStudy (anonymous):

Here's the region R: The integral is then \[\large\int_0^{\pi/3}\int_0^{\pi/2}x\cos(2x+y)\;dy\;dx\] or \[\large\int_0^{\pi/2}\int_0^{\pi/3}x\cos(2x+y)\;dx\;dy\] I think the second one is the way to go.